If the sum of reciprocal of three numbers in H.P is 18 and their product is 120 ,then find the numbers?

Grade:11

1 Answers

noogler
8 years ago
let three no. in H.P be x,y,z nd there receprocals be a,b,c which are in A.P
given a+b+c=18 and abc=120
           3b=18......b=6           because as a,b,c r in A.P  a+c=2b
    now  a+c=12 , ac=20    on solving these two eq. u get a=2, b=6, c=10
as required no. r  x,y,z  i.e,   ½  , 1/6  , 1/10
 
APPROVE IF USEFUL!!!!!!!!!!!